# An Empirical Study on Challenges for LLM Application Developers

### v3

@@ 1. Introduction
In our empirical study, we want to answer the following three research questions.
@@
Though LLM developer challenges are primarily analyzed by mining the OpenAI developer forum, our study has significant implications beyond the scope of OpenAI’s LLMs. To verify this, we reconstruct the challenge taxonomy by analyzing issues related to more LLMs (such as Llama provided by Meta AI and Gemini provided by Google) on GitHub. Therefore, our study’s findings broadly apply to LLM products developed by other companies. The commonalities in the challenges developers face (such as prompt design, the hallucination problem, and result reproducibility) are not unique to OpenAI. Still, they are prevalent across the entire LLM ecosystem. Therefore, the insights gained from our research can inform best practices, guide the development of support tools, and improve documentation and community resources for developers working with any LLM-based applications.
@@
Paper Organization. The rest of this paper is organized as follows. Section 2 introduces the background of LLM development. Section 3 describes our methods for collecting and analyzing forum posts. Section 4 analyzes the popularity of the forum. Section 5 analyzes the difficulty levels of the posts in the forum. Section 6 details our constructed taxonomy and discusses related findings. In Section 7, we first discuss the generalization of the challenge taxonomy by analyzing issues to more LLMs on GitHub. Then we analyze potential threats and the efforts to alleviate these threats. Section 8 discusses work related to our study. Section 9 summarizes our empirical study.
@@ 3. Methodology
Step 4: Constructing the taxonomy of challenges. From the 29,057 questions, we randomly selected a statistically significant sample to avoid high cost of manually analyzing all the crawled posts. This sample size ensures a 99% confidence level with a $\pm$2.5confidence interval (Aghajani et al., 2019). As a result, 2,364 questions constitute the taxonomy dataset of our study. We show the taxonomy construction methodology as follows.
@@
First, we randomly select 70% of the 2,364 posts to construct the initial taxonomy of challenges by following the previous study (Wen et al., 2021). This process involves two annotators working in a collaborative way. We use the open coding procedure (Seaman, 1999) to analyze the sampled posts, summarize the preliminary taxonomy, and construct a multi-level hierarchical taxonomy of challenges faced by OpenAI developers. Each annotator has two years of experience in OpenAI development and three years of experience in developing large language models. In this procedure, they need to carefully read the questions of the sampled posts to understand the specific challenges faced by OpenAI developers and the corresponding replies. Specifically, the open coding procedure is conducted as follows: the two annotators read all the sampled posts, with each post being read at least twice. During the analysis, they examine all the data contained in each post, including titles, problem descriptions, code snippets, and replies, summarizing the specific challenges discussed. Then they summarize the specific challenges discussed in the posts. Specifically, some posts describe issues encountered during the use of ChatGPT, for example, “Can’t log in to Chat [^6] https://community.openai.com/t/23899, which are not closely related to developers. In addition, there are also some questions related to account problems, tips for users, etc. These posts are labeled as “Unrelated”. Since the OpenAI developer forum is a platform for all OpenAI users to communicate, some posts are official announcements, document introductions, project discussions, and project sharing. For example, “I have an idea about creating a mix of emoji [^7] https://community.openai.com/t/22180, these posts are temporarily labeled as “Unrelated”. If a post is labeled as “Unrelated”, it means that these posts are not considered when constructing the taxonomy because they are not closely related to OpenAI developers. For the remaining posts, the annotators use simple phrases to summarize the developer’s challenges during the first reading. Specifically, some problem descriptions are about developers consulting on parameter settings when calling OpenAI’s APIs, for example, “GPT-4-Vision-Preview fidelity/detail-parameter [^8] https://community.openai.com/t/477563, the annotators find that this problem is asked because developers do not know how to set API parameters. Therefore, for these issues, the annotators can temporarily describe them as “API Usage Issue, Parameter Setting”.
@@
The two annotators continue to process all sampled posts according to this step and construct the taxonomy of challenges faced by OpenAI developers. The grouping process is iterative, with continuous modifications to the taxonomy based on the posts and optimization of the post descriptions to improve taxonomy quality. If a post involves multiple categories, it is assigned to all relevant categories. When the two annotators have different opinions on category determination, a third arbitrator is asked to coordinate conflicts. The third arbitrator has five years of software project development experience and three years of experience in developing large language models. After rigorous processing, all participants reach a consensus, and the challenges of all posts are classified, resulting in a preliminary taxonomy of challenges faced by OpenAI developers.
@@
Finally, we refine the taxonomy of challenges and conduct a reliability analysis. Specifically, the two annotators independently label the remaining 30% of the posts based on the generated preliminary taxonomy. They label each post as “Unrelated” or assign it to the corresponding category. If some posts can not be categorized, they are temporarily labeled as “Pending”. We use Cohen’s Kappa (k(Cohen, 1960) to measure the consistency between the two annotators during the independent labeling process. The computed $k$ value is 0.812, indicating that the two annotators almost completely agree (Landis and Koch, 1977). This demonstrates the reliability of our open coding procedure. In addition, for posts labeled as “Pending”, we introduce the third arbitrator who discusses and analyzes the posts with the two annotators to determine which category these posts ultimately belong to. If they can not be categorized into the preliminary taxonomy, new categories are added. In this phase, we incorporate five additional leaf categories into the preliminary taxonomy framework. Finally, our constructed taxonomy covers all posts related to OpenAI development, and all participants agree on the classification. Detailed result analysis can be found in Section 6.
@@ 6.1. General Questions
General questions encompass challenges that lack specific implementation details and are often posed by developers seeking foundational knowledge in LLM development. Our findings indicate that 26.3% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@ 6.2. API
This category contains challenges related to the LLM APIs. In our taxonomy, 22.9% of the total challenges fall under this category. The detailed subcategories are discussed as follows.
@@ 6.3. Generation and Understanding
Due to powerful generation and understanding capabilities, developers can use LLMs to perform a variety of tasks (such as text processing via the LLM GPT-4o, image processing via the LLM DALL-E, and audio processing via the LLM Whisper) or fine-tune the OpenAI LLMs for downstream tasks. Notice the challenges in these tasks often overlap with other challenge categories, so a question (such as parameter setting issues related to image generation [^29] https://community.openai.com/t/359438) may be classified into multiple categories. Challenges related to this category account for 19.9% of the total challenges. The detailed subcategories are discussed as follows.
@@ 6.4. Non-functional Properties
The importance of non-functional properties cannot be ignored by OpenAI developers and is closely intertwined with LLM development. These properties include aspects such as API call costs, rate limitation, and data security and privacy. In our taxonomy, 15.4% of the challenges belong to this category. The detailed subcategories are discussed as follows.
@@
Security and Privacy (D.6). Ensuring data security and preventing leakage are primary concerns for OpenAI developers, especially when using API calls that involve sensitive information. Developers should focus on safeguarding data privacy, adhering to OpenAI’s privacy policies, and securely managing API keys to avoid unauthorized access and misuse. Specifically, first, ensuring security and preventing data leakage during API calls are the main concerns for OpenAI developers. For example, developers express apprehension about potential data security risks when utilizing the API to process their uploaded files [^61] https://community.openai.com/t/354485. Additionally, they want to know whether using the API within internal networks can lead to data leakage [^62] https://community.openai.com/t/75725. Second, developers are advised to carefully review OpenAI’s privacy policy to gain insights into the data collection and processing practices during API calls and to understand the measures in place to protect users’ personal privacy. Since API keys serve as critical qualifications for developer authentication and authorization, it is imperative for developers to prioritize the secure storage and management of these keys to prevent leaks or misuse. For example, developers create a desktop application based on ChatGPT and want to know if each user can use their API key within the program for billing purposes [^63] https://community.openai.com/t/161185. Finally, developers have concerns regarding potential privacy issues associated with GPTs [^64] https://community.openai.com/t/496343. This subcategory constitutes 1.5% of all challenges.
@@ 6.5. GPT Builder
To meet specific user needs, the OpenAI organization initially introduced the ChatGPT plugin, allowing developers to design plugins with various functionalities based on user requirements. Recently, OpenAI launched GPTs with similar functionality. However, due to their later release, there are fewer posts related to GPTs. Although OpenAI has deprecated plugins and recommends using GPTs as a replacement for users’ personalized needs, the OpenAI developer forum still lists plugins as an important subforum and combines plugins and GPTs into a single subforum. Moreover, plugin development and GPT development share many similarities in terms of development goals, technical foundations, and user usage. Therefore, we call both ChatGPT plugins and GPTs “GPT Builder” and analyze the related challenges.
@@ 6.6. Prompt
Prompts play a pivotal role in LLM interactions, offering developers a mechanism to enhance the quality of the content produced. This category primarily focuses on challenges related to prompts, such as how to design and optimize them. Our findings indicate that 3.4% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@
7. Discussions
7.1. Generalization of Our Constructed Category
@@
In general, the taxonomy constructed based on the OpenAI forum covers the categories identified in the taxonomy built from GitHub. This suggests that the taxonomy developed from the OpenAI forum is representative to some extent. Through the analysis of these two taxonomies, we identify some differences between them. Specifically, we find that in the taxonomy built from GitHub, the [A.3] Feature Suggestions category accounts for 11.7%, significantly higher than the 2.9% in the OpenAI developer forum. We conjecture that since GitHub is a collaborative platform focused on software development, its users are primarily technically proficient developers who not only use existing features but also frequently contribute to the improvement of software and tools. As a result, [A.3] Feature Suggestions is a common category in GitHub Issues. The [B] API category in the GitHub challenge taxonomy also has a significantly higher proportion compared to the OpenAI developer forum (i.e., 44.7% versus 22.9%). This further indicates that GitHub users are primarily highly skilled developers who are directly involved in software development, debugging, and maintenance, and thus more frequently engage with the underlying functionality and design of APIs. For these developers, the use, optimization, and error handling of APIs are central issues in their daily development tasks, leading to a substantial amount of API-related discussions on GitHub. Finally, issues on GitHub rarely discuss topics related to [E] GPT Builder, likely because Gemini and Llama do not support user plugin or GPT development in the same way as ChatGPT. This lack of support reduces the frequency of related issues. We also notice that some categories from the taxonomy constructed based on the OpenAI forum do not appear in this taxonomy. We speculate the reasons might be: certain issues are rarely discussed in project issues (e.g., [D.4] Promotion), advanced prompt design methods have not gained sufficient attention from developers (e.g., [F.2] Retrieval Augmented Generation to [F.4] In-context Learning), and the number of projects for specific applications may be limited (e.g., [C.6] Vision Capability).

### v4

@@ 1. Introduction
In our empirical study, we want to answer the following four research questions.
@@
RQ4: Can our methodology analyze LLM developer challenges on different platforms?
Result. Though LLM developer challenges are primarily analyzed by mining the OpenAI developer forum, our study has significant implications beyond the scope of OpenAI’s LLMs. To verify this, we reconstruct the challenge taxonomy by analyzing issues related to more LLMs (such as Llama provided by Meta AI, and Gemini provided by Google) on GitHub. Therefore, our study’s findings can broadly apply to LLM products developed by other companies. The final results show that the commonalities in the challenges developers face (such as prompt design, the hallucination problem, and result reproducibility) are not unique to OpenAI. Still, they are prevalent across the entire LLM ecosystem.
In summary, the insights gained from our research can inform best practices, guide the development of support tools, and improve documentation and community resources for developers working with any LLM-based applications.
@@
  • We show the generalization of our methodology in analyzing LLM developers’ challenges by mining other platforms, such as GitHub.
@@
Paper Organization. The rest of this paper is organized as follows. Section 2 introduces the background of LLM development. Section 3 describes our methodology for collecting and analyzing forum posts. Section 4 analyzes the popularity of the forum. Section 5 analyzes the difficulty levels of the posts in the forum. Section 6 details our constructed taxonomy and discusses related findings. Section 7 discusses the generalization of our methodology by analyzing issues from more LLMs on GitHub. Section 8 first reviews and synthesizes the key findings and offers actionable implications. Then we analyze potential threats and the efforts to alleviate these threats. Section 9 discusses work related to our study. Section 10 summarizes our empirical study.
@@ 3. Methodology
Step 4: Constructing the taxonomy of challenges. From the 29,057 questions, we randomly selected a statistically significant sample to avoid the high cost of manually analyzing all the crawled posts. This sample size ensures a 99% confidence level with a $\pm$2.5confidence interval (Aghajani et al., 2019). As a result, 2,364 questions constitute the taxonomy dataset of our study. We show the taxonomy construction methodology as follows.
@@
First, we randomly select 70% of the 2,364 posts to construct the initial taxonomy of challenges by following the previous study (Wen et al., 2021). This process involves two annotators working in a collaborative way. We use the open coding procedure (Seaman, 1999) to analyze the sampled posts, summarize the preliminary taxonomy, and construct a multi-level hierarchical taxonomy of challenges faced by LLM developers. Each annotator has two years of experience in OpenAI development and three years of experience in developing large language models. In this procedure, they need to carefully read the questions of the sampled posts to understand the specific challenges faced by LLM developers and the corresponding replies. Specifically, the open coding procedure is conducted as follows: the two annotators read all the sampled posts, with each post being read at least twice. During the analysis, they examine all the data contained in each post, including titles, problem descriptions, code snippets, and replies, summarizing the specific challenges discussed. Then they summarize the specific challenges discussed in the posts. Specifically, some posts describe issues encountered during the use of ChatGPT, for example, “Can’t log in to Chat [^6] https://community.openai.com/t/23899, which are not closely related to developers. In addition, there are also some questions related to account problems, tips for users, etc. These posts are labeled as “Unrelated”. Since the OpenAI developer forum is a platform for all OpenAI users to communicate, some posts are official announcements, document introductions, project discussions, and project sharing. For example, “I have an idea about creating a mix of emoji [^7] https://community.openai.com/t/22180, these posts are temporarily labeled as “Unrelated”. If a post is labeled as “Unrelated”, it means that these posts are not considered when constructing the taxonomy because they are not closely related to LLM developers. For the remaining posts, the annotators use simple phrases to summarize the developer’s challenges during the first reading. Specifically, some problem descriptions are about developers consulting on parameter settings when calling OpenAI’s APIs, for example, “GPT-4-Vision-Preview fidelity/detail-parameter [^8] https://community.openai.com/t/477563, the annotators find that this problem is asked because developers do not know how to set API parameters. Therefore, for these issues, the annotators can temporarily describe them as “API Usage Issue, Parameter Setting”.
@@
The two annotators continue to process all sampled posts according to this step and construct the taxonomy of challenges faced by LLM developers. The grouping process is iterative, with continuous modifications to the taxonomy based on the posts and optimization of the post descriptions to improve taxonomy quality. If a post involves multiple categories, it is assigned to all relevant categories. When the two annotators have different opinions on category determination, a third arbitrator is asked to coordinate conflicts. The third arbitrator has five years of software project development experience and three years of experience in developing large language models. After rigorous processing, all participants reach a consensus, and the challenges of all posts are classified, resulting in a preliminary taxonomy of challenges faced by LLM developers.
@@
Finally, we refine the taxonomy of challenges and conduct a reliability analysis. Specifically, the two annotators independently label the remaining 30% of the posts based on the generated preliminary taxonomy. They label each post as “Unrelated” or assign it to the corresponding category. If some posts can not be categorized, they are temporarily labeled as “Pending”. We use Cohen’s Kappa (k(Cohen, 1960) to measure the consistency between the two annotators during the independent labeling process. The computed $k$ value is 0.812, indicating that the two annotators almost completely agree (Landis and Koch, 1977). This demonstrates the reliability of our open coding procedure. In addition, for posts labeled as “Pending”, we introduce the third arbitrator who discusses and analyzes the posts with the two annotators to determine which category these posts ultimately belong to. If they can not be categorized into the preliminary taxonomy, new categories are added. In this phase, we incorporate five additional leaf categories into the preliminary taxonomy framework. Finally, our constructed taxonomy covers all posts related to OpenAI development, and all participants agree on the classification. Detailed result analysis can be found in Section 6. Finally, we also apply this methodology to analyze the challenges of LLM application developers by mining GitHub issues related to more LLMs and the detailed results can be found in Section 7.
@@ 6.1. General Questions
General Questions encompass challenges that lack specific implementation details and are often posed by developers seeking foundational knowledge in LLM development. These challenges reflect the diverse issues developers face when starting with or working on LLM-based applications. Common difficulties include understanding how to integrate LLMs with custom applications, addressing conceptual gaps (e.g., token usage and API functionalities), and suggesting enhancements to existing features. Our findings indicate that 26.3% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@ 6.2. API
This category contains challenges related to the LLM APIs, including issues with fault handling, error messages, and parameter configurations, which collectively hinder development efficiency and create barriers to effective usage. Common challenges include managing unsatisfactory outputs, troubleshooting error messages, and understanding or optimizing API parameters for specific tasks. These challenges highlight the complexity of integrating LLM APIs into practical applications and underscore the need for clearer documentation and support mechanisms. In our taxonomy, 22.9% of the total challenges fall under this category. The detailed subcategories are discussed as follows.
@@ 6.3. Generation and Understanding
Due to powerful generation and understanding capabilities, developers can use LLMs to perform a variety of tasks (such as text processing via the LLM GPT-4o, image processing via the LLM DALL-E, and audio processing via the LLM Whisper) or fine-tune the OpenAI LLMs for downstream tasks. However, developers frequently face challenges in effectively leveraging these capabilities. Common issues include troubleshooting errors in text, image, audio, and embedding processing tasks; managing fine-tuning processes; and addressing API-related limitations. These challenges often involve optimizing input prompts, resolving parameter configuration problems, and addressing unexpected API errors or quality inconsistencies in outputs. Notice the challenges in these tasks often overlap with other challenge categories, so a question (such as parameter setting issues related to image generation [^29] https://community.openai.com/t/359438) may be classified into multiple categories. Challenges related to this category account for 19.9% of the total challenges. The detailed subcategories are discussed as follows.
@@ 6.4. Non-functional Properties
The importance of non-functional properties cannot be ignored by LLM developers and is closely intertwined with LLM development. These properties include aspects such as API call costs, rate limitations, and data security and privacy, which significantly influence the efficiency and usability of LLM-based applications. Developers frequently face challenges such as managing API call costs by optimizing token usage, understanding and working within rate limitations, ensuring compliance with regulatory requirements, handling token limitations for large-scale datasets, and safeguarding data security and privacy during API usage. In our taxonomy, 15.4% of the challenges belong to this category. The detailed subcategories are discussed as follows.
@@
Security and Privacy (D.6). Ensuring data security and preventing leakage are primary concerns for LLM developers, especially when using API calls that involve sensitive information. Developers should focus on safeguarding data privacy, adhering to OpenAI’s privacy policies, and securely managing API keys to avoid unauthorized access and misuse. Specifically, first, ensuring security and preventing data leakage during API calls are the main concerns for LLM developers. For example, developers express apprehension about potential data security risks when utilizing the API to process their uploaded files [^61] https://community.openai.com/t/354485. Additionally, they want to know whether using the API within internal networks can lead to data leakage [^62] https://community.openai.com/t/75725. Second, developers are advised to carefully review OpenAI’s privacy policy to gain insights into the data collection and processing practices during API calls and to understand the measures in place to protect users’ personal privacy. Since API keys serve as critical qualifications for developer authentication and authorization, it is imperative for developers to prioritize the secure storage and management of these keys to prevent leaks or misuse. For example, developers create a desktop application based on ChatGPT and want to know if each user can use their API key within the program for billing purposes [^63] https://community.openai.com/t/161185. Finally, developers have concerns regarding potential privacy issues associated with GPTs [^64] https://community.openai.com/t/496343. This subcategory constitutes 1.5% of all challenges.
@@ 6.5. GPT Builder
To meet specific user needs, the OpenAI organization initially introduced the ChatGPT plugin, allowing developers to design plugins with various functionalities based on user requirements. Recently, OpenAI launched GPTs with similar functionality. However, due to their later release, there are fewer posts related to GPTs. Although OpenAI has deprecated plugins and recommends using GPTs as a replacement for users’ personalized needs, the OpenAI developer forum still lists plugins as an important subforum and combines plugins and GPTs into a single subforum. Moreover, plugin development and GPT development share many similarities in terms of development goals, technical foundations, and user usage. Therefore, we call both ChatGPT plugins and GPTs “GPT Builder” and analyze the related challenges. LLM developers commonly face challenges during the development and testing phases of GPT Builders. These include selecting suitable development environments and tools, addressing technical errors (such as server or parsing issues), and managing the complexity of integrating multiple GPTs for specific tasks. During testing, challenges often arise from validation errors, incorrect functionality, and debugging failures. Our findings indicate that 12.1% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@ 6.6. Prompt
Prompts play a pivotal role in LLM interactions, offering developers a mechanism to enhance the quality of the content produced. This category primarily focuses on challenges related to prompts, such as how to design and optimize them effectively. Developers frequently face issues in crafting prompts that achieve specific goals, modifying prompts to enhance output quality, and leveraging advanced techniques like retrieval-augmented generation, chain of thought, and in-context learning. Our findings indicate that 3.4% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@
7. RQ4: Generalization of Our Methodology
@@
In general, the taxonomy constructed based on the OpenAI forum covers the categories identified in the taxonomy built from GitHub. This suggests that the taxonomy developed from the OpenAI forum is representative to some extent. However, our analysis reveals some differences between the two taxonomies, which provide insights into the unique characteristics of these platforms and their user bases.
First, the [A.3] Feature Suggestions category accounts for 11.7% in the GitHub taxonomy, significantly higher than the 2.9% observed in the OpenAI developer forum. This disparity highlights the collaborative nature of GitHub as a platform where technically proficient developers not only use existing features but actively contribute to the improvement and evolution of software and tools. These developers often engage in discussions about potential enhancements and new functionalities for tools, reflecting a proactive role in shaping the development landscape. In contrast, the OpenAI developer forum serves a broader audience, including beginners and normal users, where discussions are often centered around usage challenges rather than feature proposals.
Second, the [B] API category constitutes 44.7% of the GitHub taxonomy compared to 22.9% in the OpenAI developer forum. This significant difference underscores the emphasis GitHub users place on API-related issues, including usage, optimization, and error handling. This focus is indicative of a user base that is deeply involved in software development and debugging processes, where APIs are integral to their daily tasks. The prevalence of API discussions on GitHub reflects the critical role APIs play in collaborative software projects and the need for highly technical solutions.
Third, issues related to [E] GPT Builder are rarely discussed in the GitHub taxonomy, whereas they represent a notable category in the OpenAI developer forum. This discrepancy can be attributed to the fact that LLMs like Gemini and Llama do not support user plugins or GPT development in the same way as ChatGPT, leading to a lower frequency of GPT Builder-related discussions on GitHub.
Finally, some categories identified in the OpenAI forum taxonomy are absent from the GitHub taxonomy. For instance, [D.4] Promotion is rarely discussed on GitHub, likely because the platform is centered around collaborative coding rather than promoting individual tools or plugins. Similarly, advanced prompt design techniques such as [F.2] Retrieval Augmented Generation and [F.4] In-context Learning are underrepresented, suggesting that these methods have not yet gained widespread attention among developers on GitHub. Finally, categories like [C.6] Vision Capability may be limited due to the relatively small number of projects focused on such specific applications on GitHub.
These differences highlight the unique characteristics and developer priorities on each platform. GitHub caters to a more technical and collaborative audience, leading to a focus on API optimization and feature suggestions. In contrast, the OpenAI forum accommodates a diverse range of users, resulting in broader discussions that include usage challenges and advanced techniques.
8. Discussions
In this section, we first review and synthesize the key findings, offer implications from these findings, and highlight opportunities for future work. We second analyze potential threats to our empirical study.
8.1. Findings and Implications of Our Empirical Study
This study analyzed the challenges faced by LLM application developers, constructing a taxonomy based on forum discussions to capture the breadth of these challenges. Our key findings can be summarized as follows:
  • Diverse Developer Challenges. The taxonomy reveals a wide range of challenges developers face, spanning functional and non-functional aspects of LLM development. Key categories include issues with API usage, GPT Builder development, prompt design, and advanced techniques.
  • Significance of API-Related Issues. API-related challenges dominate the taxonomy, reflecting their central role in developer workflows. These include technical difficulties such as error handling, parameter configuration, and performance optimization, underscoring the need for robust support and documentation.
  • Emerging Interest in GPT Builders. The presence of GPT Builder challenges highlights the growing interest in customizing LLM functionality. Developers frequently encounter issues during the development and testing phases, suggesting a need for more targeted resources in this area.
  • Underutilization of Advanced Prompt Techniques. Despite their potential to enhance LLM outputs, advanced techniques like CoT and RAG remain underutilized, indicating gaps in developer awareness and accessibility.
  • Non-Functional Concerns Are Pervasive. Challenges related to cost management, rate limitations, and data privacy reflect the importance of non-functional properties in LLM usage. These issues often act as barriers to adoption, especially for developers working with limited resources.
Based on these findings, we provide the following key actionable implications for LLM-related stakeholders.
  • Enhanced Documentation and Resources. To address the diverse challenges identified, LLM vendors should prioritize the development of comprehensive documentation, including tutorials, best practices, and sample configurations. Specific focus should be given to API optimization, prompt engineering, and advanced techniques like CoT and RAG.
  • Tooling for Non-Functional Properties. Non-functional concerns such as cost and rate limitations necessitate tools for real-time tracking and optimization. Developing related tools for monitoring usage and predicting costs could significantly improve developer experiences.
  • Supporting GPT Builder Development. The prevalence of GPT Builder challenges underscores the need for streamlined development and testing tools. LLM vendors could introduce IDEs or automated testing frameworks tailored for GPT customization.
  • Promoting Advanced Techniques. The limited use of advanced techniques suggests a need for targeted education initiatives. LLM vendors and the broader software engineering community should provide workshops, case studies, and automated prompt design tools to promote their adoption.